sqrt(x+1)+1 = sqrt(2x)
Answer by mananth(16946)


sqrt(x+1)+1 = sqrt(2x)
sqrt(x+1)= sqrt(2x)-1

square both sides
x+1 = 2x+2sqrt(2x)+1
x-2x= 2sqrt(2x)
-x= 2sqrt(2x)
square
x^2=4*2x
x^2=8x
x=8